Rose’s garden is in the shape of a trapezoid. If the height of the trapezoid is 16m, one base is 20m, and the area is 224 square meters, find the length of the other base
Answer by checkley71(8403)

16(20+X)/2=224
(320+16X)/2=224 CROSS MULTIPLY
320+16X=448
16X=448-320
16X=128
X=128/16
X=8 IS THE LENGTH OF THE OTHER SIDE.
PROOF
16(20+8)/2=224
(16*28)/2=224
448/2=224
224=224